A 2-digit number , its unit digit is twice its tens digit. If the product of the two digits equals the third of the number.What is the number ?
Answer by Boreal(15235)

xy are the digits
x*y=(1/3)(10x+y); the 10x comes from the fact that the x is in the tens column
but y=2x
2x^2=(1/3)(10x+2x)=4x
2x^2-4x=0
2x(x-2)=0
x=2; (x=0 does not make sense)
y=4
number is 24
The product of the digits (8) = 1/3 of the number (24).